# Quantifying task-relevant representational similarity using decision variable correlation

### v3

@@ 1 Introduction
Deep learning [bengioRepresentationLearningReview2014, lecunDeepLearning2015] has substantially impacted how neuroscientists construct brain models. For vision neuroscience, deep neural networks offer candidate models for studying the primate ventral pathway [yaminsPerformanceoptimizedHierarchicalModels2014a, khaligh-razaviDeepSupervisedNot2014, yaminsUsingGoaldrivenDeep2016] and, more recently, dorsal pathway [ebrahimpourVentralDorsalNeuralNetworks2020, mineaultYourHeadThere2021, choiDualStreamNeuralNetwork2023]. Early work reported that the representations in convolutional neural networks (CNNs) trained on image categorization tasks can explain a substantial fraction of variance in high-level visual areas, surpassing classic models for these areas [yaminsPerformanceoptimizedHierarchicalModels2014a]. Follow-up research has tested many variants of deep networks on their alignment with the brain using both neural data [muttenthalerHumanAlignmentNeural2023, wangIncorporatingNaturalLanguage2022, conwellWhatCan182023] and behavior data [rajalinghamLargeScaleHighResolutionComparison2018, geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021]. One appealing hypothesis is that deep networks that exhibit higher accuracy and robustness in vision tasks, or trained on larger datasets would better explain visual processing in the brain.
@@
One important question is how to compare deep network models and the brain. One class of methods seeks to quantify the similarity of internal representations between models and brains. This includes methods such as representational similarity analysis (RSA) [kriegeskorteRepresentationalSimilarityAnalysis2008a], linear regression [yaminsUsingGoaldrivenDeep2016, naselarisEncodingDecodingFMRI2011], and generalized shape metrics [williamsGeneralizedShapeMetrics2022a, Harvey2024representational]. More recently, another class of methods that put more emphasis on quantifying the behavioral similarity has been proposed, including Cohen’s Kappa [geirhosAccuracyQuantifyingTrialbytrial2020] and I2n behavioral predictivity [rajalinghamLargeScaleHighResolutionComparison2018, schrimpfBrainScoreWhichArtificial2020]. The goal of these methods is to provide an image-by-image comparison of the decision strategies used by neural networks and the brain. One challenge has been how to properly disentangle the model accuracy, decision biases and decision consistency from behavior [sebastianDecisionvariableCorrelation2018, geirhosAccuracyQuantifyingTrialbytrial2020]. Intriguingly, studies seem to find contradictory trends in model-brain alignment. While some studies suggest that deep learning systems converge to learning a common representation [Huh2024platonic, Shen2025alignment], others suggest that the similarity will not grow indefinitely or has capped [Fel2022harmonizing, Linsley2025better]. The reason for this remains unclear.
@@
Here, we propose a principled approach that combines the merits of model comparisons at the representation and behavior levels. Our method is based on decision variable correlation (DVC) developed to measure the behavioral similarity based on choice data [sebastianDecisionvariableCorrelation2018] in signal detection theory, and we have generalized it to analyze the consistency of high-dimensional neural representations. While prior work developed techniques to estimate DVCs from behavioral data [sebastianDecisionvariableCorrelation2018], we instead infer DVCs from neural representations. Thus, our proposed DVC metric should be interpreted as a measure of the representational similarity. Our approach specifically quantifies the trial-by-trial consistency of two neural representations for solving a classification task, ignoring features that are irrelevant for the task. In doing so, our approach enables one to infer the consistency of the decision strategy of two observers from their internal representations based on the assumption of optimal linear readouts.
@@ 2 Background and relevant work
Model architecture and scale One hypothesis has been that as models improve in task performance or architectural complexity, their internal representations become more brain-like. The Brain-Scores [schrimpfBrainScoreWhichArtificial2020] of the image classification models were reported to be positively correlated to ImageNet-1k accuracy, although the trend plateaus at higher accuracy. On the other hand, studies using RSA reported that neither model scale nor architecture significantly improved alignment to human behavioral similarity judgments [muttenthalerHumanAlignmentNeural2023]. Another study using RSA reported negative correlation between alignment to human neural activity and model complexity [sartzetakiOneHundredNeural2024]. Studies using Cohen’s Kappa reported that human-model behavioral consistency at the image level remains low despite improved performance on out-of-distribution datasets with scaling[geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021].
@@
Robustness The primate visual system is robust against external and internal noise, prompting the question of whether robustness to adversarial perturbations or corruptions is related to brain-model alignment [dapelloAligningModelMacaque2022]. Recent work proposed that adversarial robustness might promote the learning of representations better aligned with human perception [engstromAdversarialRobustnessPrior2019]. By enforcing alignment with monkey IT representations, models exhibited both enhanced adversarial robustness and increased behavioral alignment with human subjects [dapelloAligningModelMacaque2022]. Another study found that model metamers – artificial stimuli that elicit the same response as natural stimuli, generated by robust models– are more recognizable to humans, but are not themselves predictive of recognizability [featherModelMetamersReveal2023]. However, studies using Cohen’s Kappa report that robust models still diverge from humans in their error patterns[geirhosPartialSuccessClosing2021]. Additionally, there has also been evidence suggesting that even though adversarial training increased model robustness, these robust networks may not use human-like features unless explicitly aligned [Subramanian2023spatial, Linsley2023adversarial].
@@
Rich and multimodal training data Using Cohen’s Kappa, [geirhosPartialSuccessClosing2021] reports that models trained on larger and more diverse datasets become more human-like in their behaviors. On the other hand, a recent large-scale study using a variation of RSA reported that upgrading from ImageNet-1k to ImageNet21k does not significantly improve alignment to human brain, but object-oriented ImageNet datasets lead to much better alignment than datasets containing only places or faces [conwellWhatCan182023]. Similarly, an ecologically-motivated dataset seems to improve model-brain alignment [mehrerEcologicallyMotivatedImage2021]. Joint vision-language models such as CLIP have also been shown to better predict human brain activity [geirhosPartialSuccessClosing2021, wangIncorporatingNaturalLanguage2022, subramaniamRevealingVisionLanguageIntegration2024].
@@
Similarity measures Classic methods such as linear predictivity measure how well neural responses can be predicted from network representation [yaminsPerformanceoptimizedHierarchicalModels2014a]. Representational similarity analysis (RSA) [kriegeskorteRepresentationalSimilarityAnalysis2008a, khaligh-razaviDeepSupervisedNot2014] is a popular approach that quantifies the similarity in the geometrical structure of two representations, and is blind to the specific features used to solve a task [dujmovicInferringDNNBrainAlignment2024]. Other shape metrics such as centered kernel alignment (CKA) and canonical correlation analysis (CCA) also do not specifically address the similarity in task-relevant features. Measures of behavioral similarity such as error consistency are task-focused [geirhosAccuracyQuantifyingTrialbytrial2020], yet they may be overcalibrated on accuracies of the observers and are therefore sensitive to the choice of decoders[geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021] as we will show later. Recent studies highlighted the challenges in the interpretation of results based on these methods [schrimpfBrainScoreWhichArtificial2020, conwellWhatCan182023], e.g., different methods for quantifying the brain-model similarity could lead to different conclusions [soniConclusionsNeuralNetwork2024, sucholutskyGettingAlignedRepresentational2023]. As the field continues to develop methods for analyzing and interpreting model-brain alignments [Barbosa2025quantifying, Chou2025feature], it would be desirable to have principled, task-relevant, accuracy-agnoistic measures to better illustrate the possible divergence between brains and models.
@@ 3 DVC: Quantifying the trial-by-trial consistency of two representations
We develop a new method to evaluate the consistency of two neural representations. This method is based on a principled generalization of signal detection theory. It enables one to estimate how correlated the decision strategies of two observers are for a classification task. The method is robust to the observers’ biases and is not confounded by the behavioral accuracy. It operates at the level of neural representations, and enables one to analyze the internal representation to infer the consistency of the two representations for solving the classification tasks. Thus, the method can quantify task-relevant representational similarity. Compared to methods purely based on behavior [sebastianDecisionvariableCorrelation2018, geirhosAccuracyQuantifyingTrialbytrial2020], it takes advantage of the richness of the internal representations of neural networks and brains. Meanwhile, in contrast to methods for analyzing the similarity of two neural representations (such as representational similarity analysis), our method focuses on the dimensions that are relevant for a behavioral task and is invariant to variability along other task-irrelevant dimensions.
@@ 3.1 Decision variable correlations (DVC) of two neural representations
Signal detection theory is fundamental in the study of perceptual behavior. The idea is that, for binary-choice tasks, observer uses a continuous decision variable (DV) to make a choice (Fig. 1a). Recently, [sebastianDecisionvariableCorrelation2018] proposed to generalize signal detection theory to study the correlation of decision variables of two observers (Fig. 1b). Their method inferred the DVC from binary choice data. Here, we develop a simple new strategy to infer trial-by-trial DVCs from high-dimensional internal representations (Fig. 1c).
@@ 3.2 Implementation of the DVC method
We next discuss how we implement the DVC framework to analyze the high-dimensional neural representations from the brains or deep networks.
@@ 4.1 High trial-by-trial consistency of V4 & IT representations across monkey brains
We evaluate the consistency of neural representations in different macaque monkeys. We used the publicly available dataset of objects rendered on naturalistic scenes [majajSimpleLearnedWeighted2015]. In these experiments, they used images from eight classes {animals, boats, cars, chairs, faces, fruits, planes, tables}, with 400 images each, totaling 400 $\times$ 8 = 3200 images. Recordings were taken from V4 and IT areas of two adult macaque monkeys passively viewing these images. The brain representation is taken to be the time-binned spike counts averaged over  50 repeats. 100 neurons from V4 and IT combined were obtained from each monkey.
@@ 4.2 Deep networks with higher accuracy on ImageNet exhibit lower DVCs with brains
We study a set of models (n=43, obtained from Torchvision, [marcelTorchvisionMachinevisionPackage2010]) pretrained on ImageNet-1k, an influential benchmark in image classification. This also offers a fair comparison between models by controlling for confounding factors related to different training data. We use the same neural datasets from [majajSimpleLearnedWeighted2015] as above. We feed the images in [majajSimpleLearnedWeighted2015] into deep vision-based neural networks, subject to standard transforms. The model representation for an image is defined as the activation taken from the penultimate layer – the last layer before the final logit layer.
@@
Network vs. network We next examine the DVCs between different deep neural networks. Specifically, we evaluate DVCs between deep networks from different model families [^3] We define a model family as a set of architectures sharing a canonical computational graph – such as residual, attention, or convolutional block structures—with variation limited to hyperparameters like depth, width, patch size, or token embedding dimension. See Appendix B for more model details. Using DVC, models from the same family or otherwise share architectural similarities are clearly clustered together(Fig. 2a,b). We find that DVCs between pairs of models within the same family (similar model structures and training processes) are substantially higher than pairs from different families (p = 1.33e-56, Fig. 2d), consistent with previous findings [kornblithSimilarityNeuralNetwork2019]. We also find DVCs between models not to be exceedingly high. Despite being trained on the same dataset, they do not seem to converge to a single solution, at least not significantly higher compared to the similarity between the two monkeys (Fig. 2d). These results imply that model structures and training processes still play significant roles in the task solutions found by the models. Notably, these results differ substantially from results obtained by computing error consistency. These studies [geirhosAccuracyQuantifyingTrialbytrial2020, geirhosPartialSuccessClosing2021] reported that (i) the consistency between model and brain is very low; (ii) the consistency between network models is much higher than the consistency between humans. Later, we will address the difference between the methodologies.
@@ 4.3 Adversarially trained networks, while highly consistent, have low DVCs to the brain
Robustness represents one important difference between deep networks and our perceptual systems. Small perturbations to images that are imperceptible to humans can lead to qualitative errors in deep networks (i.e., adversarial examples) [szegedyIntriguingPropertiesNeural2014a]. Adversarial examples reflect the misalignment between representations in deep networks and brains, given certain local perturbations in the inputs. Adversarial robustness can be increased by using adversarial training, e.g., by finding adversarial examples and adding them to the training set. Studies suggest that features learned through adversarial training may be more aligned with human perception [engstromAdversarialRobustnessPrior2019, featherModelMetamersReveal2023], posing an intriguing hypothesis that by making networks locally consistent with human perception, network representations may be better aligned with brain representations globally.
@@ 4.4 Networks trained on rich datasets exhibit no increase in DVC to the brain
Whereas ImageNet-1k has been an important benchmark dataset for the image classification community for a decade, recent state-of-the-art models are trained on larger datasets such as ImageNet-21k, which is a scaled-up version of ImageNet-1k, and JFT-300M, which is proprietary. Models trained on larger, more diverse datasets may generalize within a larger domain, and may show better out-of-distribution generalization ability. A recent study showed that models trained on these larger datasets may exhibit better alignment with human behavior [geirhosPartialSuccessClosing2021]. Furthermore, the negative correlations between classification performance and DVCs to the brains (Fig. 2c) suggest the possibility of overfitting to a particular dataset that is much smaller than what brains are trained on evolutionarily, developmentally, and during the experiments [kaznatcheevNothingMakesSense2022]. Therefore, it is of particular interest to investigate whether models trained on the richer datasets exhibit higher DVCs to the brain.
@@
Figure 5: Comparsion to Cohen’s Kappa. (a) Heatmaps showing the DVC and Cohen’s Kappa for pairs of representative models. (b) There is a strong positive correlation between Cohen’s Kappa and DVC on model-model consistency (evaluated on this dataset) (c) There is a decent positive correlation between Cohen’s Kappa and DVC on model-monkey consistency. (d) The values of Cohen’s Kappa between models (blue) and human subjects (green) are low while Cohen’s Kappa between models and between human subjects are high, consistent with the original report. Re-analyzed based on data from [geirhosImageNettrainedCNNsAre2018]. (e) Scatter plot showing the relationship between Cohen’s Kappa and the difference in accuracy of pairs of observers, and the theoretical upper bound. (f) The response histogram to ’edges’ distortion based on the model and decision rules used in [geirhosAccuracyQuantifyingTrialbytrial2020] and the original study [geirhosImageNettrainedCNNsAre2018]. Different colors represent different nerual networks. (g) Simulation results show that Cohen’s Kappa is sensitive to decision biases, while DVC is invariant to decision biases.

### v4

@@ 1 Introduction
Deep learning [3, 26] has substantially impacted how neuroscientists construct brain models. For vision neuroscience, deep neural networks offer candidate models for studying the primate ventral pathway [53, 22, 52] and, more recently, dorsal pathway [12, 32, 4]. Early work reported that the representations in convolutional neural networks (CNNs) trained on image categorization tasks can explain a substantial fraction of variance in high-level visual areas, surpassing classic models for these areas [53]. Follow-up research has tested many variants of deep networks on their alignment with the brain using both neural data [33, 47, 7] and behavior data [36, 16, 17]. One appealing hypothesis is that deep networks that exhibit higher accuracy and robustness in vision tasks, or trained on larger datasets would better explain visual processing in the brain.
@@
One important question is how to compare deep network models and the brain. One class of methods seeks to quantify the similarity of internal representations between models and brains. This includes methods such as representational similarity analysis (RSA) [25], linear regression [52, 34], and generalized shape metrics [49, 19]. More recently, another class of methods that put more emphasis on quantifying the behavioral similarity has been proposed, including Cohen’s Kappa [16] and I2n behavioral predictivity [36, 38]. The goal of these methods is to provide an image-by-image comparison of the decision strategies used by neural networks and the brain. One challenge has been how to properly disentangle the model accuracy, decision biases and decision consistency from behavior [39, 16]. Intriguingly, studies seem to find contradictory trends in model-brain alignment. While some studies suggest that deep learning systems converge to learning a common representation [20, 40], others suggest that the similarity will not grow indefinitely or has capped [15, 28]. The reason for this remains unclear.
@@
Here, we propose a principled approach that combines the merits of model comparisons at the representation and behavior levels. Our method is based on decision variable correlation (DVC) developed to measure the behavioral similarity based on choice data [39] in signal detection theory, and we have generalized it to analyze the consistency of high-dimensional neural representations. While prior work developed techniques to estimate DVCs from behavioral data [39], we instead infer DVCs from neural representations. Thus, our proposed DVC metric should be interpreted as a measure of the representational similarity. Our approach specifically quantifies the trial-by-trial consistency of two neural representations for solving a classification task, ignoring features that are irrelevant for the task. In doing so, our approach enables one to infer the consistency of the decision strategy of two observers from their internal representations based on the assumption of optimal linear readouts.
@@ 2 Background and relevant work
Model architecture and scale One hypothesis has been that as models improve in task performance or architectural complexity, their internal representations become more brain-like. The Brain-Scores [38] of the image classification models were reported to be positively correlated to ImageNet-1k accuracy, although the trend plateaus at higher accuracy. On the other hand, studies using RSA reported that neither model scale nor architecture significantly improved alignment to human behavioral similarity judgments [33]. Another study using RSA reported negative correlation between alignment to human neural activity and model complexity [37]. Studies using Cohen’s Kappa reported that human-model behavioral consistency at the image level remains low despite improved performance on out-of-distribution datasets with scaling[16, 17].
@@
Robustness The primate visual system is robust against external and internal noise, prompting the question of whether robustness to adversarial perturbations or corruptions is related to brain-model alignment [9]. Recent work proposed that adversarial robustness might promote the learning of representations better aligned with human perception [13]. By enforcing alignment with monkey IT representations, models exhibited both enhanced adversarial robustness and increased behavioral alignment with human subjects [9]. Another study found that model metamers – artificial stimuli that elicit the same response as natural stimuli, generated by robust models– are more recognizable to humans, but are not themselves predictive of recognizability [14]. However, studies using Cohen’s Kappa report that robust models still diverge from humans in their error patterns[17]. Additionally, there has also been evidence suggesting that even though adversarial training increased model robustness, these robust networks may not use human-like features unless explicitly aligned [44, 27].
@@
Rich and multimodal training data Using Cohen’s Kappa, [17] reports that models trained on larger and more diverse datasets become more human-like in their behaviors. On the other hand, a recent large-scale study using a variation of RSA reported that upgrading from ImageNet-1k to ImageNet21k does not significantly improve alignment to human brain, but object-oriented ImageNet datasets lead to much better alignment than datasets containing only places or faces [7]. Similarly, an ecologically-motivated dataset seems to improve model-brain alignment [31]. Joint vision-language models such as CLIP have also been shown to better predict human brain activity [17, 47, 43].
@@
Similarity measures Classic methods such as linear predictivity measure how well neural responses can be predicted from network representation [53]. Representational similarity analysis (RSA) [25, 22] is a popular approach that quantifies the similarity in the geometrical structure of two representations, and is blind to the specific features used to solve a task [11]. Other shape metrics such as centered kernel alignment (CKA) and canonical correlation analysis (CCA) also do not specifically address the similarity in task-relevant features. Measures of behavioral similarity such as error consistency are task-focused [16], yet they may be overcalibrated on accuracies of the observers and are therefore sensitive to the choice of decoders[16, 17] as we will show later. Recent studies highlighted the challenges in the interpretation of results based on these methods [38, 7], e.g., different methods for quantifying the brain-model similarity could lead to different conclusions [42, 45]. As the field continues to develop methods for analyzing and interpreting model-brain alignments [1, 5], it would be desirable to have principled, task-relevant, accuracy-agnoistic measures to better illustrate the possible divergence between brains and models.
@@ 3 DVC: Quantifying the trial-by-trial consistency of two representations
We develop a new method to evaluate the consistency of two neural representations. This method is based on a principled generalization of signal detection theory. It enables one to estimate how correlated the decision strategies of two observers are for a classification task. The method is robust to the observers’ biases and is not confounded by the behavioral accuracy. It operates at the level of neural representations, and enables one to analyze the internal representation to infer the consistency of the two representations for solving the classification tasks. Thus, the method can quantify task-relevant representational similarity. Compared to methods purely based on behavior [39, 16], it takes advantage of the richness of the internal representations of neural networks and brains. Meanwhile, in contrast to methods for analyzing the similarity of two neural representations (such as representational similarity analysis), our method focuses on the dimensions that are relevant for a behavioral task and is invariant to variability along other task-irrelevant dimensions.
@@ 3.1 Decision variable correlations (DVC) of two neural representations
Signal detection theory is fundamental in the study of perceptual behavior. The idea is that, for binary-choice tasks, observer uses a continuous decision variable (DV) to make a choice (Fig. 1a). Recently, [39] proposed to generalize signal detection theory to study the correlation of decision variables of two observers (Fig. 1b). Their method inferred the DVC from binary choice data. Here, we develop a simple new strategy to infer trial-by-trial DVCs from high-dimensional internal representations (Fig. 1c).
@@ 3.2 Implementation of the DVC method
We next discuss how we implement the DVC framework to analyze the high-dimensional neural representations. The code is available at https://github.com/wei-bbc-lab/DVC.
@@ 4.1 High trial-by-trial consistency of V4 & IT representations across monkey brains
We evaluate the consistency of neural representations in different macaque monkeys. We used the publicly available dataset of objects rendered on naturalistic scenes [29]. In these experiments, they used images from eight classes {animals, boats, cars, chairs, faces, fruits, planes, tables}, with 400 images each, totaling 400 $\times$ 8 = 3200 images. Recordings were taken from V4 and IT areas of two adult macaque monkeys passively viewing these images. The brain representation is taken to be the time-binned spike counts averaged over  50 repeats. 100 neurons from V4 and IT combined were obtained from each monkey.
@@ 4.2 Deep networks with higher accuracy on ImageNet exhibit lower DVCs with brains
We study a set of models (n=43, obtained from Torchvision, [30]) pretrained on ImageNet-1k, an influential benchmark in image classification. This also offers a fair comparison between models by controlling for confounding factors related to different training data. We use the same neural datasets from [29] as above. We feed the images in [29] into deep vision-based neural networks, subject to standard transforms. The model representation for an image is defined as the activation taken from the penultimate layer – the last layer before the final logit layer.
@@
Network vs. network We next examine the DVCs between different deep neural networks. Specifically, we evaluate DVCs between deep networks from different model families [^3] We define a model family as a set of architectures sharing a canonical computational graph – such as residual, attention, or convolutional block structures—with variation limited to hyperparameters like depth, width, patch size, or token embedding dimension. See Appendix B for more model details. Using DVC, models from the same family or otherwise share architectural similarities are clearly clustered together(Fig. 2a,b). We find that DVCs between pairs of models within the same family (similar model structures and training processes) are substantially higher than pairs from different families (p = 1.33e-56, Fig. 2d), consistent with previous findings [24]. We also find DVCs between models not to be exceedingly high. Despite being trained on the same dataset, they do not seem to converge to a single solution, at least not significantly higher compared to the similarity between the two monkeys (Fig. 2d). These results imply that model structures and training processes still play significant roles in the task solutions found by the models. Notably, these results differ substantially from results obtained by computing error consistency. These studies [16, 17] reported that (i) the consistency between model and brain is very low; (ii) the consistency between network models is much higher than the consistency between humans. Later, we will address the difference between the methodologies.
@@ 4.3 Adversarially trained networks, while highly consistent, have low DVCs to the brain
Robustness represents one important difference between deep networks and our perceptual systems. Small perturbations to images that are imperceptible to humans can lead to qualitative errors in deep networks (i.e., adversarial examples) [46]. Adversarial examples reflect the misalignment between representations in deep networks and brains, given certain local perturbations in the inputs. Adversarial robustness can be increased by using adversarial training, e.g., by finding adversarial examples and adding them to the training set. Studies suggest that features learned through adversarial training may be more aligned with human perception [13, 14], posing an intriguing hypothesis that by making networks locally consistent with human perception, network representations may be better aligned with brain representations globally.
@@ 4.4 Networks trained on rich datasets exhibit no increase in DVC to the brain
Whereas ImageNet-1k has been an important benchmark dataset for the image classification community for a decade, recent state-of-the-art models are trained on larger datasets such as ImageNet-21k, which is a scaled-up version of ImageNet-1k, and JFT-300M, which is proprietary. Models trained on larger, more diverse datasets may generalize within a larger domain, and may show better out-of-distribution generalization ability. A recent study showed that models trained on these larger datasets may exhibit better alignment with human behavior [17]. Furthermore, the negative correlations between classification performance and DVCs to the brains (Fig. 2c) suggest the possibility of overfitting to a particular dataset that is much smaller than what brains are trained on evolutionarily, developmentally, and during the experiments [21]. Therefore, it is of particular interest to investigate whether models trained on the richer datasets exhibit higher DVCs to the brain.
@@
Figure 5: Comparsion to Cohen’s Kappa. (a) Heatmaps showing the DVC and Cohen’s Kappa for pairs of representative models. (b) There is a strong positive correlation between Cohen’s Kappa and DVC on model-model consistency (evaluated on this dataset) (c) There is a decent positive correlation between Cohen’s Kappa and DVC on model-monkey consistency. (d) The values of Cohen’s Kappa between models (blue) and human subjects (green) are low while Cohen’s Kappa between models and between human subjects are high, consistent with the original report. Re-analyzed based on data from [18]. (e) Scatter plot showing the relationship between Cohen’s Kappa and the difference in accuracy of pairs of observers, and the theoretical upper bound. (f) The response histogram to ’edges’ distortion based on the model and decision rules used in [16] and the original study [18]. Different colors represent different nerual networks. (g) Simulation results show that Cohen’s Kappa is sensitive to decision biases, while DVC is invariant to decision biases.